Pickett Industries just paid a dividend of D0 = $1.10. Analysts expect the company's dividend to grow by 20% this year, by 10% in Year 2, and at a constant rate of 3% in Year 3 and thereafter. The required return on this low-risk stock is 8.00%. What is the per-share estimate of the stock’s intrinsic value?
Answer #1

Last Dividend, D0 = $1.10

Growth rate for this year is 20%, next year is 10% and a constant growth rate (g) is 3%

Required Return, rs = 8.00%

D1 = $1.10 * 1.20 = $1.32
D2 = $1.32 * 1.10 = $1.452
D3 = $1.452 * 1.03 = $1.49556

P2 = D3 / (rs - g)
P2 = $1.49556 / (0.08 - 0.03)
P2 = $29.9112

P0 = $1.32/1.08 + $1.452/1.08^2 + $29.9112/1.08^2
P0 = $28.11

So, the per-share estimate of the stock’s intrinsic value is $28.11